# Page Time of Primordial Black Holes in the Standard Model and Beyond

### v2

@@ III Entropy production and Page time
We computed ${\cal N}_{ilm}$ for 150 values of $x$ in the range $[0.001,3.0]$ and for 100 values of $a_{\star}$ in the range $[0,0.9999]$. For the $l,m$ modes, the spectra were obtained for $m\in[-l,l]$, with $l$ values considered up to 9 for bosons and up to $19/2$ for fermions. To validate our numerical computation, we obtained the following values for each spin particle per internal degree of freedom for the entropy production function $\zeta_{i}$, assuming massless degrees of freedom and a Schwarzschild BH
$\zeta_{i}(M,0)=10^{-3}\times\begin{cases}3.4648&\text{for }s_{i}=0,\\ 1.6855&\text{for }s_{i}=1/2,\\ 0.6338&\text{for }s_{i}=1,\\ 0.0651&\text{for }s_{i}=2.\end{cases}$ (18)
We find a $0.00551$‰, $-0.580$‰, and $2.17$‰ difference for $s_{i}=1/2,1,2$, respectively, with the values presented in Refs. [129, 11, 12].
@@
determines by the amount of entropy in the emitted radiation times the decrease of the Bekenstein-Hawking entropy. For BHs heavier than $M\gtrsim 10^{17}~{}{\rm g}$, emitting only neutrinos, and massless photons and gravitons, we find that $\beta=1.61862$, a value with a difference of $-48.66$ parts-per-million (ppm) with the value given in Refs. [129, 11, 12]. Neutrino emission is expected to become suppressed when $M\gtrsim 10^{26}~{}{\rm g}$, assuming the lightest neutrino to have a mass of $0.01~{}{\rm eV}$ [68]. Thus, for heavier BHs, the entropy production ratio drops to $\beta=1.48471$, with a difference with the value given in Ref. [12] of $6.6$ ppm. For BH masses below $M\lesssim 10^{17}~{}{\rm g}$, additional SM and potentially beyond-the-SM degrees of freedom are emitted, altering the $\beta$ ratio from the values discussed earlier. As an example, for a Schwarzschild BH with an initial mass of $10^{9}~{}{\rm g}$, emitting all SM degrees of freedom but no additional particles, we find $\beta=1.64064$.
@@
[…]
@@ IV.1 Phenomenological Consequences
Figure 6 presents the percentage of particles emitted before the Page time as a function of the initial spin parameter $a_{\star}^{\rm in}$ for a PBH with an initial mass of $M_{\rm in}=10^{9}~{}\rm g$, assuming the emission of SM degrees of freedom plus an additional massless particle, which can be a scalar (green), fermion (purple), vector (light blue), or spin-2 particle (red). We find that for Schwarzschild BHs, approximately $37.8\%$ of the total emission occurs before the Page time, regardless of the particle type. This result can be understood by noting that the evaporation rate for massless particles remains constant throughout the lifetime of a Schwarzschild BH, even though the absolute emission rate depends on the specific particle type. Therefore, in this scenario, the contribution to $\Delta N_{\rm eff}$ will be impacted if either the PBH lifetime or the particle production rates are altered after the Page time.
The situation differs for Kerr BHs. As discussed earlier, the emission of higher-spin particles is enhanced for larger values of $a_{\star}^{\rm in}$. Consequently, for gravitons, we observe that approximately $99\%$ of the total graviton production occurs before the Page time for $a_{\star}^{\rm in}\to 1$. However, for lower spin values, this percentage decreases, reaching only about $60\%$ for $a_{\star}^{\rm in}\sim 0.4$. A similar trend is observed for vector particles, though their pre-Page-time emission fraction reaches only about $70\%$ for $a_{\star}^{\rm in}\to 1$. In contrast, the emission behavior of fermions and scalars exhibits an opposite trend. For highly spinning BHs, the fraction of emission occurring before the Page time is reduced, with only $\sim 18\%$ of scalars and $\sim 36\%$ of fermions emitted for a nearly maximally rotating PBH.
@@
Fig. 6: Percentage of particles emitted before the Page time for a Kerr PBH with an initial mass of $M_{\rm in}=10^{9}~{}\rm g$, assuming the emission of all SM degrees of freedom along with a single massless, non-interacting particle. The additional particle is considered to be either a scalar (green), fermion (purple), vector (light blue), or graviton (red).

### v3

@@ III Entropy production and Page time
We computed ${\cal N}_{ilm}$ for 150 values of $x$ in the range $[0.001,3.0]$ and for 100 values of $a_{\star}$ in the range $[0,0.9999]$. For the $l,m$ modes, the spectra were obtained for $m\in[-l,l]$, with $l$ values considered up to 9 for bosons and up to $19/2$ for spin 1/2 fermions. For Rarita–Schwinger fermions, we followed the procedure in Ref. [130] to compute ${\cal N}_{ilm}$ up to a value of $l=11/2$. To validate our numerical computation, we obtained the following values for each spin particle per internal degree of freedom for the entropy production function $\zeta_{i}$, assuming massless degrees of freedom and a Schwarzschild BH
$\zeta_{i}(M,0)=10^{-3}\times\begin{cases}3.4648&\text{for }s_{i}=0,\\ 1.6855&\text{for }s_{i}=1/2,\\ 0.6338&\text{for }s_{i}=1,\\ 0.2110&\text{for }s_{i}=3/2,\\ 0.0651&\text{for }s_{i}=2.\end{cases}$ (18)
We find a 0.005510.00551‰, 0.580-0.580‰, and 2.172.17‰ difference for $s_{i}=1/2,1,2$, respectively, with the values presented in Refs. [129, 11, 12].
@@
determines by the amount of entropy in the emitted radiation times the decrease of the Bekenstein-Hawking entropy. For BHs heavier than $M\gtrsim 10^{17}~{\rm g}$, emitting only neutrinos, and massless photons and gravitons, we find that $\beta=1.61862$, a value with a difference of 48.66-48.66 parts-per-million (ppm) with the value given in Refs. [129, 11, 12]. Neutrino emission is expected to become suppressed when $M\gtrsim 10^{26}~{\rm g}$, assuming the lightest neutrino to have a mass of $0.01~{\rm eV}$ [68]. Thus, for heavier BHs, the entropy production ratio drops to $\beta=1.48471$, with a difference with the value given in Ref. [12] of $6.6$ ppm. For BH masses below $M\lesssim 10^{17}~{\rm g}$, additional SM and potentially beyond-the-SM degrees of freedom are emitted, altering the $\beta$ ratio from the values discussed earlier. As an example, for a Schwarzschild BH with an initial mass of $10^{9}~{\rm g}$, emitting all SM degrees of freedom but no additional particles, we find $\beta=1.64064$.
@@
Now, turning to Kerr BHs, Fig. 2 illustrates the entropy ratio $\beta$ for a BH with a mass of $M=10^{9}~{\rm g}$ as a function of the initial spin parameter $a_{\star}^{\rm in}$. The top panel displays the dimensionless quantities corresponding to the change in Bekenstein-Hawking entropy, $GMdS_{\tt BH}/dt$, (blue curve) and the von Neumann entropy associated with the emitted radiation, $GMdS_{\rm rad}/dt$, (dashed red curve), while the bottom panel shows the $\beta$ parameter. A notable distinction from the Schwarzschild case is that, for a BH emitting all SM degrees of freedom, the thermodynamic entropy change rate becomes zero at $a_{\star}\approx 0.9215$ [^2] This value depends on the set of particles emitted, which is determined by the instantaneous BH temperature. Ref. [132] reports a value of $a_{\star}\approx 0.8868$ for a BH emitting two massless neutrinos, photons, and gravitons.. Above this threshold, the rate of change becomes positive, signifying an increase in entropy. This behavior arises because, for $a_{\star}\gtrsim 0.9215$, the emission of superradiant modes dominates over thermal emission, which lead to an increase in entropy, as discussed in Ref. [132]. The $\beta$ parameter captures this transition, becoming negative for $a_{\star}\gtrsim 0.9215$ and turning positive once the BH spin parameter falls below such a value.
@@ IV.1 Phenomenological Consequences
Figure 6 presents the percentage of particles emitted before the Page time as a function of the initial spin parameter $a_{\star}^{\rm in}$ for a PBH with an initial mass of $M_{\rm in}=10^{9}~\rm g$, assuming the emission of SM degrees of freedom plus an additional massless particle, which can be a scalar (green), fermion (purple), vector (light blue), spin 3/2 (yellow), or spin-2 particle (red). We find that for Schwarzschild BHs, approximately $37.8\%$ of the total emission occurs before the Page time, regardless of the particle type. This result can be understood by noting that the evaporation rate for massless particles remains constant throughout the lifetime of a Schwarzschild BH, even though the absolute emission rate depends on the specific particle type. Therefore, in this scenario, the contribution to $\Delta N_{\rm eff}$ will be impacted if either the PBH lifetime or the particle production rates are altered after the Page time.
The situation differs for Kerr BHs. As discussed earlier, the emission of higher-spin particles is enhanced for larger values of $a_{\star}^{\rm in}$. Consequently, for gravitons, we observe that approximately $99\%$ of the total graviton production occurs before the Page time for $a_{\star}^{\rm in}\to 1$. However, for lower spin values, this percentage decreases, reaching only about $60\%$ for $a_{\star}^{\rm in}\sim 0.4$. A similar trend is observed for vector and spin 3/2 particles, though their pre-Page-time emission fraction reaches only about $70\%$ and $90\%$ for $a_{\star}^{\rm in}\to 1$, respectively. In contrast, the emission behavior of fermions and scalars exhibits an opposite trend. For highly spinning BHs, the fraction of emission occurring before the Page time is reduced, with only $\sim 18\%$ of scalars and $\sim 36\%$ of fermions emitted for a nearly maximally rotating PBH.
@@
Fig. 6: Percentage of particles emitted before the Page time for a Kerr PBH with an initial mass of $M_{\rm in}=10^{9}~\rm g$, assuming the emission of all SM degrees of freedom along with a single massless, non-interacting particle. The additional particle is considered to be either a scalar (green), fermion (purple), vector (light blue), spin 3/2 (yellow), or graviton (red).